Evaluate ∫0^π\2α α sin αx dx,α>0 by using Leibniz formula.


1
Expert's answer
2021-08-16T02:19:24-0400

Compute the definite integral:

"\int_0^{\frac{1}{2}\pi \alpha} \alpha \sin (\alpha x)dx\\"

Factor out constants:

"\alpha \int_0^{\frac{1}{2}\pi \alpha} \sin (\alpha x)dx\\"

For the integrand "\sin \alpha x" , substitute "u= \alpha x" and "du = \alpha dx" . This gives a new lower bound "u=0 \space \alpha = 0" and upper bound "u= \alpha (0.5 \pi \alpha)= 0.5 \pi \alpha^2"

"\int_0^{\frac{1}{2}\pi \alpha^2} \sin (u)du\\"

Apply the fundamental theorem of calculus.

The antiderivative of "sin(u) \space is -cos(u)"

"=(-\cos (u))|_{0}^{0.5 \pi \alpha^2}"

Evaluate the antiderivative at the limits and subtract.

"=(-\cos (u))|_{0}^{0.5 \pi \alpha^2}= (-\cos (0.5 \pi \alpha^2))-(-\cos (0))\\
=1-\cos(0.5 \pi \alpha^2)"